OpenStudy (anonymous):

1/2 + radical 3/ 2 (x-pi/6)

OpenStudy (anonymous):

I think you meant to type \[\frac{1}{2}+\sqrt[3]{\left(x-\frac{\pi}{6}\right)^2 }\] yes?

OpenStudy (anonymous):

If so, what is your question?